If tan (a+b-c) / tan (a-b+c) = tan c / tan b


 


What is the value of sin 2a + sin 2b + sin 2c.


 

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear karthik,

Multiply through by cos(a-b+c).cos(b).cos(a-b+c).cos(c) …

sin(a+b-c)cos(a-b+c).sin(b).cos(c) − sin(a-b+c).cos(a+b-c).sin(c).cos(b) = 0

½.sin(b).cos(c){ sin(2a) + sin(2b-2c) } − ½.sin(c).cos(b){ sin(2a) − sin(2b-2c) } = 0

½.sin(2a).sin(b-c) + ½.sin(2b-2c){ sin(b).cos(c) + sin(c).cos(b) } = 0

½.sin(2a).sin(b-c) + sin(b-c).cos(b-c).sin(b+c) = 0

½.sin(2a).sin(b-c) + ½.sin(b-c){ sin(2b) + sin(2c) } = 0

½.sin(b-c){ sin(2a) + sin(2b) + sin(2c) } = 0

∴ sin(2a) + sin(2b) + sin(2c) = 0

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi